This afternoon we got it going with two more track repairs on the motor drives and another on the front panel.
It is now showing a mono picture.
A badly dented crystal has emerged from the cabinet and we have not yet found out where it has come from or even if it actually belongs to this set. Adrian has taken the crystal away with him so there is no photo right now.
I have not done a decoder on one of these machines as most of my trade experience comes from industrial CCTV stuff so we now need an expert on colour faults on these kind of sets.
We now have a running deck all be it a bit rough from a set that was wrecked by a courier.
The take up spool is a little sluggish starting up but that is most likely a bad capacitor downstream of the power supply so Adrian can manage that himself.
Adrian has another example of this set and it has HIFI audio trouble and this one is showing good audio although we did not connect it up as my test monitor only has mono sound and we would have had to run long cables to a big heavy valve stereo amp over the other side of the room.
This job could end up with an audio board borrow to check it out and see if it is heads or electronics on the second set.
I have not looked at the second set yet as this one was mostly courier damage and squeaky hubs.
Chris has had some success today too so we are doing all right